Arrange the following molecules in order of increasing average molecular speed. NH3 at 450K, Ar at 311K, NH3 at 367K, N2 at 367K Enter formulas and temperatures in the boxes below: 1 = slowest, 4 = fastest
Answer #1

Answer: -

Ar at 311 K ----- 1 ----- slowest

N2 at 367 K ----- 2  

NH3 at 367 K ----- 3   

NH3 at 450 K ----- 4 ----- fastest

-----------------------------------------------------------------------------------------

Formula: -

Average molecular speed (V) of a gas is calculated as

   ...........(1)

Where,

V = Average molecular speed of a gas.

R = Gas constant = 8.314 J/mol.K

T = Absolute Temperature in kevin

M = Molar mass of gas

----------------------------------------------------------------------------------------------------

1) NH3  at 450 K

for NH3 M = 17 g/mol = kg/mol = 0.017 kg/mol

T = 450 K

From equation (1) we have,

i.e.   

but, we know,

1 J/kg = 1 m2/s2

therefore,

----------------------------------------------------------------------------------------------------------

2) Ar at 311 K

for Ar M = 39.9 g/mol = kg/mol = 0.0399 kg/mol

T = 311 K

From equation (1) we have,

i.e.   

but, we know,

1 J/kg = 1 m2/s2

therefore,

--------------------------------------------------------------------------------------------------------

3) NH3  at 367 K

for NH3 M = 17 g/mol = kg/mol = 0.017 kg/mol

T = 367 K

From equation (1) we have,

i.e.   

but, we know,

1 J/kg = 1 m2/s2

therefore,

-------------------------------------------------------------------------------------------------------------

4) N2 at 367 K

for N2 M = 28 g/mol =    kg/mol = 0.028 kg/mol

T = 367 K

From equation (1) we have,

i.e.   

but, we know,

1 J/kg = 1 m2/s2

therefore,

-------------------------------------------------------------------------------------------------------------